OpenStudy (anonymous):

i need help .. can you teach me how to get the checking of this problem ? 2x²-3x-9=0 i know the answer of that but i don't know how to check if it is right or wrong please help me :D

OpenStudy (anonymous):

substitute in the origin equation \[2x ^{2}-3x-9=??\] if the answer is zero then you are right

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{orange}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (yanasidlinskiy):

Tell us what you got:D

OpenStudy (anonymous):

2x^2 - 3x -9 1 -3 2 3 cross multiply 1*3=3 and 2 * -3 = -6 then add 3 - 6 = -3 well its correct the answer will be (x - 3)(2x + 3) x= 3 and -3/2

hero (hero):

If you have found the x values, plug them in to the equation to make sure you get zero for both values.